OpenStudy (anonymous):

Find f(5) for this sequence: f(1) = 2 and f(2) = 4, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______

OpenStudy (gorv):

f(3)=f(1)+f(2)+f(2)=2+4+4=10 f(4)=f(1)+f(2)+f(3)=2+4+10=16 f(5)=f(1)+f(2)+f(4)=2+4+16=22